OpenStudy (anonymous):

Doug's house is behind a baseball field and he wants to place a sign on the corner of his fence that tells the distance (in feet) that the sign is from home plate.

OpenStudy (anonymous):

Doug's house is behind a baseball field and he wants to place a sign on the corner of his fence that tells the distance (in feet) that the sign is from home plate. image that shows the location of home plate, left field, and the corner of the fence He cannot directly measure the distance from home plate to the corner of the fence due to a retention pond that is between the corner of the fence and home plate. Instead, he measures the distance from the corner of his fence to the spot just behind the left field foul pole and finds it is 552 feet. He estimates the angle from the corner of the fence to home plate at 36 degrees. Using the cosine ratio, he gets the following equation: cos36º = 552/x 0.890 = 552/x 0.890x = 552 x = 682 The distance from home Instructions: Find the length of an inaccessible distance by using a similar method. , draw a picture of your situation.

OpenStudy (anonymous):

 tan 36 = y / 552 y = 552 * tan 36 = 401 feet

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

welcome